# Graph Anomaly Detection via Adaptive Test-time Representation Learning across Out-of-Distribution Domains

### v1

@@ 4.1 Backbone GNN and Projection head
The backbone architecture of AdaGraph-T3 is a Graph Neural Network (GNN) that operates on source and target graph data. Our model can apply any GNN architecture, including Message-passing Neural Networks (MPNNs). We denote the set of neighbors at node $v$ as $\mathcal{N}(v)=\{u|(u,v)\in\mathcal{E}\}$. For each layer $\ell\in\{1,2,\ldots L\}$ of the GNN the information from neighboring nodes of $v$ is aggregated via message-passing as follows:
@@
where $M(\cdot)$ is an element-wise and permutation-invariant operator (e.g., average, max, or min). The embedding of node $v$ at the $\ell$-th layer is derived from the aggregated neighbors’ embeddings and the previous embedding of $v$:
@@
where $\sigma(\cdot)$ is a non-linear activation function, ||||| | denotes the concatenation operator, $W^{\ell}\in\mathbb{R}^{p_{\ell+1}\times 2p_{\ell}}$, and $\mathbf{b}^{\ell}\in\mathbb{R}^{p_{\ell+1}}$ are trainable weight matrices and bias vectors. Representations from the last layer ($\ell=L$) can be converted to binary class probabilities using the $\mathrm{sigmoid}$ function.
To generate initial representations $\mathbf{h}_{v}^{0}$ for nodes with different feature sets in $D_{s}$ and $D_{t}$, GATD3 learns projection matrices (or encoders) $P_{s}$ and $P_{t}$ as follows:
$\mathbf{h}_{v}^{0}=P_{s}.\mathbf{x}_{v},$ (3)
where $P_{s}\in\mathbb{R}^{p\times p_{s}}$ represents a trainable projection for the source data and $\mathbf{x}_{v}$ are node features. The same is performed for target data using a projection matrix $P_{t}$. These encoders enable projecting source and target data into a shared representation space not only for transfer learning but also for interpreting target predictions as we will show in the experiments.
Without loss of generality, we apply GraphSAGE [10] as AdaGraph-T3’s backbone MPNN architecture. GraphSAGE applies node sampling in the aggregation phase to increase its scalability.
@@ 4.2 Homophily-based Self-supervised Loss
Test-time training requires a self-supervised loss $\mathcal{L}_{self}$ applied during both training and test-time training stages. We propose the use of a previously introduced homophily-based affinity score, which is tailored for unsupervised anomaly detection [26, 4].
The local affinity score measures the similarity or connection strength between a node and its neighbors. It is based on the one-class homophily phenomenon, where normal nodes tend to have a stronger affinity with their neighbors compared with anomalies. This enables these scores to be applied for unsupervised GAD. To the best of our knowledge, our work is the first to apply affinity scores for the GAD problem within the test-time training framework.
More formally, the local affinity score is computed by comparing a node’s representation against those of its immediate neighbors in the graph using various similarity metrics such as cosine similarity, Euclidean distance, Jensen-Shannon divergence, and Wasserstein distance [20, 5, 28]. By taking the average similarity with neighbors of a node $v$, we obtain a single anomaly score $s(v)$ based on how well a node is associated with its local graph structure:
@@
where $N(v_{i})$ denotes the neighbors of node $v_{i}$, $\mathbf{h}_{u}^{L}$ represents the learned node embedding for $u$, and $\mathrm{sim}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}^{T}\mathbf{b}}{\lVert\mathbf{a}\rVert\lVert\mathbf{b}\rVert}$ describes the cosine similarity. The unsupervised task consists of maximizing the affinity score for each node.
@@
$\mathcal{L}_{\text{self}}=s({v})+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (5)
where $\mathcal{L}_{\text{reg}}=\sum_{i\in\mathcal{V}}\left(\frac{\sum_{j\in\mathcal{N}_{i}^{-}}\text{sim}(h_{i},h_{j})}{|\mathcal{N}_{i}^{-}|}\right)$ minimizes the similarity between non-connected nodes to maintain structural distinctiveness in the embedding space, with $\mathcal{N}_{i}^{-}$ representing the set of nodes not connected to node $i$ and $\lambda$ controlling the regularization strength.
It has been noted by previous work [26, 4] that enforcing similarity across all connected nodes might inadvertently cause non-homophily nodes to become similar. Normal Structure-preserved Graph Truncation (NSGT) attempts to preserve normal node structure by making binary decisions to remove edges between dissimilar nodes based on distance. However, this approach risks losing important structural information and requires setting explicit thresholds for edge removal decisions. In the next section, we introduce our enhanced message-passing approach through attention-based NSAW, a simpler and more flexible approach to suppress the adversarial effect of non-homophily edges in equation 5.
@@ 4.3 Normal Structure-preserved Attention Weighting (NSAW)
We propose Normal Structure-preserved Attention Weighting (NSAW) as an alternative to Normal Structure-preserved Graph Truncation [26]. NSAW applies a robust extension to graph attention [32] as a more flexible way to suppress the effect of anomalous nodes on the representation of normal nodes during message-passing. These attention weights are learned end-to-end as part of the GNN training process.
@@
NSAW’s symmetric construction provides a crucial defensive mechanism through two complementary aspects. First, when a normal node $v$ connects to both normal nodes $\{v_{1},v_{2},\ldots,v_{k}\}$ and anomalous nodes $\{a_{1},a_{2},\ldots,a_{m}\}$, the attention naturally assigns high weights to normal neighbors and low weights to anomalous ones due to feature similarity. As illustrated in the Figure 3, when a normal node assigns low attention $\alpha(v\rightarrow a)\approx 0$ to an anomalous node (thin arrow), even if the anomalous node attempts to assign high attention back (thick arrow), the minimum operation then ensures that the reverse influence is also minimized as $\alpha(a\rightarrow v)=\min(\alpha(a\rightarrow v),\alpha(v\rightarrow a))\approx 0$ (shown by the dashed arrow), regardless of the original attention weight. Second, through the learnable attention parameters $U^{l}$ and the self-supervised loss, the architecture gradually learns to suppress the influence of anomalous patterns during training, further strengthening this defensive mechanism through the GNN’s message passing structure.
Figure 3: Illustration of the symmetric attention mechanism. Node $a$ assigns a high attention weight $\alpha_{1}$ to node $v$, while node $v$ assigns a lower attention weight $\alpha_{2}$ to node $a$. Using the minimum operation, the final symmetric attention weight becomes $\alpha_{2}$.
The computed attention weights modulate the message passing in the GNN architecture. For a node $v$, the attention-weighted messages from its neighbors are aggregated as follows:
$\mathbf{h}_{v}^{\ell}\!=\!M(\{\widetilde{A}^{\ell-1}[u,v]\cdot\mathbf{h}_{u}^{\ell-1}\!\mid\!u\!\in\mathcal{N}(v)\})$ (8)
@@ 4.4 Class-aware Regularization
$\mathcal{L}_{s}=\sum_{i\in\mathcal{V}}\left(\frac{\sum_{j\in\mathcal{N}_{i}^{-}}w_{j}\cdot\text{sim}(h_{i},h_{j})}{|\mathcal{N}_{i}^{-}|}\right)$ (9)
@@
where $\alpha>1$ is the weighting factor for anomalous nodes. We set $\alpha$ inversely proportional to the class ratio This regularization can only be applied during source training where we have access to node labels, while target training relies solely on the homophily-based loss due to its unsupervised nature.
@@ 4.5 Test-Time Training
We can now describe our framework for anomaly detection with domain adaptation (AdaGraph-T3), summarized in Figure 2. Following the test-time training paradigm, AdaGraph-T3 is based on two tasks, the main (supervised) task and the auxiliary self-supervised learning (SSL) task. In the training phase, AdaGraph-T3 is trained based on both tasks using a joint loss minimized based on the source dataset. Subsequently, the test-time training phase leverages only the SSL task using the target dataset, which is assumed to be unlabeled.
Training phase. During training phase, AdaGraph-T3 applies the source graph $\mathcal{G}_{s}$ to learn an encoder $\theta_{s}$, a decoder $\theta_{m}$, and a predictor $\theta_{\text{pred}}$ by minimizing a loss function $\mathcal{L}_{\text{train}}$ that is a combination of a supervised loss $\mathcal{L}_{\text{sup}}$, for which we apply cross-entropy, and a self-supervised loss $\mathcal{L}_{\text{self}}$, for which we apply the homophily-based affinity scores introduced in Section 4.2:
$\mathcal{L}_{\text{sup}}=\sum_{v\in\mathcal{V}_{s}}y_{v}\log{\hat{y}_{v}}+\lambda_{s}\mathcal{L}_{s}$ (11)
$\mathcal{L}_{\text{self}}=\frac{1}{|N(v)|}\sum_{u\in N(v)}\mathrm{sim}(\mathbf{h}_{v}^{L},\mathbf{h}_{u}^{L})+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}$ (12)
$\mathcal{L}_{\text{train}}=\sum_{v\in\mathcal{V}_{s}}\Bigl{(}\mathcal{L}_{\text{sup}}(v,y;\theta_{m},\theta_{s},\theta_{\text{pred}})+\lambda\mathcal{L}_{\text{self}}(v;\theta_{m},\theta_{s})\Bigr{)}$ (13)
@@
We describe the source encoder $\theta_{s}$ (projection) and decoder $\theta_{m}$ (GNN) in Section 4.1. The predictor $\theta_{\text{pred}}$ is a multi-layer perception that maps node embeddings to label predictions.
@@
$\mathcal{L}_{\text{TTT}}=\sum_{v\in\mathcal{V}_{t}}\lambda_{t}\mathcal{L}_{\text{self}}(v;\theta_{t},\theta_{m})$ (14)
@@
where $\mathcal{L}_{\text{self}}$ is defined in Equation 12 and the GNN decoder $\theta_{m}$ is re-used from the training phase to embed target nodes.
Inference. After the TTT phase, anomalies are detected in the target dataset by combining the target encoder $\theta_{t}$ learned during the TTT phase with the GNN decoder $\theta_{m}$ and the predictor $\theta_{\text{pred}}$, both learned during the training phase. Intuitively, the source $\theta_{s}$ and target $\theta_{t}$ encoders project features from the source and target graph, respectively, into a new shared feature space. The GNN $\theta_{m}$ operates over this shared space to generate node embeddings where anomalies can be distinguished from normal nodes using the predictor $\theta_{\text{pred}}$.

### v2

@@ 4.1 Backbone GNN and Projection head
The backbone architecture of GADT3 is a Graph Neural Network (GNN) that operates on source and target graph data. Our model can apply any GNN architecture, including Message-passing Neural Networks (MPNNs). We denote the set of neighbors at node $v$ as $\mathcal{N}(v)=\{u|(u,v)\in\mathcal{E}\}$. For each layer $\ell\in\{1,2,\ldots L\}$ of the GNN the information from neighboring nodes of $v$ is aggregated via message-passing as follows:
@@
where $M(\cdot)$ is an element-wise and permutation-invariant operator. The embedding of node $v$ at the $\ell$-th layer is derived from the aggregated neighbors’ embeddings and the previous embedding of $v$:
@@
where $\sigma(\cdot)$ is a non-linear activation function, $||$ denotes the concatenation operator, $W^{\ell}\in\mathbb{R}^{p_{\ell+1}\times 2p_{\ell}}$, and $\mathbf{b}^{\ell}\in\mathbb{R}^{p_{\ell+1}}$ are trainable weight matrices and bias vectors. Representations from the last layer ($\ell=L$) can be converted to binary class probabilities using the $\mathrm{sigmoid}$ function. To generate initial representations $\mathbf{h}_{v}^{0}=P_{s}.\mathbf{x}_{v}$ for nodes with different feature sets in $D_{s}$ and $D_{t}$, GADT3 learns projection matrices (or encoders) $P_{s}$ and $P_{t}$ where, $P_{s}\in\mathbb{R}^{p\times p_{s}}$ is a trainable projection for the source data and $\mathbf{x}_{v}$ are node features. The same is performed for target data using the matrix $P_{t}$. These encoders enable projecting source and target data into a shared representation space, facilitating transfer learning.
Without loss of generality, we apply GraphSAGE (Hamilton et al., 2017) as GADT3’s backbone MPNN architecture. GraphSAGE applies node sampling in the aggregation phase to increase its scalability.
@@ 4.2 Homophily-based Self-supervised Loss
Test-time training requires a self-supervised loss $\mathcal{L}_{self}$ applied during both training and test-time training stages. We propose the use of a previously introduced homophily-based affinity score, which is tailored for unsupervised anomaly detection (Qiao & Pang, 2024; Chen et al., 2024b).
The local affinity score measures the similarity or connection strength between a node and its neighbors. It is based on the one-class homophily phenomenon, where normal nodes tend to have a stronger affinity with their neighbors compared with anomalies. This enables these scores to be applied to unsupervised GAD. To the best of our knowledge, our work is the first to apply affinity scores for the GAD problem within the test-time training framework.
More formally, the local affinity score is computed by comparing a node’s representation against those of its immediate neighbors in the graph using various similarity metrics such as cosine similarity, Euclidean distance, Jensen-Shannon divergence, and Wasserstein distance (Lin, 1991; Chen & Liu, 2020; Shen et al., 2018). By taking the average similarity with neighbors of a node $v$, we obtain a single anomaly score $s(v)$ based on how well a node is associated with its local graph structure:
@@
where $N(v_{i})$ denotes the neighbors of $v_{i}$, $\mathbf{h}_{u}^{L}$ represents the learned embedding for $u$, and $\mathrm{sim}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}^{T}\mathbf{b}}{\lVert\mathbf{a}\rVert\lVert\mathbf{b}\rVert}$ describes the cosine similarity. The unsupervised task consists of maximizing the affinity score for each node.
@@
$\mathcal{L}_{\text{self}}=-\sum_{v\in\mathcal{V}}s({v})+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (4)
where $\mathcal{L}_{\text{reg}}=\sum_{v\in\mathcal{V}}\left(\frac{\sum_{u\in\mathcal{N}_{v}^{-}}\text{sim}(\mathbf{h}_{v},\mathbf{h}_{u})}{|\mathcal{N}_{v}^{-}|}\right)$ minimizes the similarity between non-connected nodes to maintain structural distinctiveness in the embedding space, with $\mathcal{N}_{v}^{-}$ representing the set of nodes not connected to node $v$ and $\lambda$ controlling the regularization strength.
This homophily loss also serves as an implicit alignment mechanism, encouraging local consistency among neighboring node embeddings in the target graph. Since node labels are unavailable during training, the homophily loss is applied to all connected node pairs. This may include anomalous pairs, but under our homophily assumption, most edges still connect similar (normal) nodes, making the signal reliable in practice.
By leveraging the frozen source decoder $\theta_{m}$, the model preserves the source-learned graph patterns, achieving alignment without requiring explicit alignment terms or target labels.
It has been noted by previous work (Qiao & Pang, 2024; Chen et al., 2024b) that enforcing similarity across all connected nodes might inadvertently cause non-homophily nodes to become similar. Normal Structure-preserved Graph Truncation (NSGT) attempts to preserve normal node structure by making binary decisions to remove edges between dissimilar nodes based on distance. However, this approach risks losing important structural information and requires setting explicit thresholds for edge removal decisions. In the next section, we introduce our enhanced message-passing approach through attention-based NSAW, a simpler and more flexible approach to suppress the adversarial effect of non-homophily edges in equation 4.
@@ 4.3 Normal Structure-preserved Attention Weighting (NSAW)
We propose Normal Structure-preserved Attention Weighting (NSAW) as an alternative to Normal Structure-preserved Graph Truncation (Qiao & Pang, 2024). NSAW applies a robust extension to graph attention (Veličković et al., 2017) as a more flexible way to suppress the effect of anomalous nodes on the representation of normal nodes during message-passing. Specifically, by downweighting abnormal-to-normal connections, NSAW reduces the influence of structurally inconsistent neighbors. This preserves the structure of normal regions and encourages more distinct representations for anomalies. These attention weights are learned end-to-end as part of the GNN training process.
@@
Figure 3: Symmetric attention: node $a$ gives high attention to node $v$ ($\alpha_{1}$), while $v$ gives lower attention to $a$ ($\alpha_{2}$). The final symmetric weight is $\min(\alpha_{1},\alpha_{2})=\alpha_{2}$.
NSAW’s symmetric construction provides a crucial defensive mechanism through two complementary aspects. First, when a normal node $v$ connects to both normal nodes $\{v_{1},v_{2},\ldots,v_{k}\}$ and anomalous nodes $\{a_{1},a_{2},\ldots,a_{m}\}$, the attention naturally assigns high weights to normal neighbors and low weights to anomalous ones due to feature similarity. As illustrated in the Figure 3, when a normal node assigns low attention $\alpha(v\rightarrow a)\approx 0$ to an anomalous node (thin arrow), even if the anomalous node attempts to assign high attention back (thick arrow), the minimum operation ensures that the reverse influence is also minimized as $\alpha(a\rightarrow v)=\min(\alpha(a\rightarrow v),\alpha(v\rightarrow a))\approx 0$ (dashed arrow), regardless of the original attention weight. Second, through the learnable attention parameters $U^{l}$ and the self-supervised loss, the architecture gradually learns to suppress the influence of anomalous patterns during training, further strengthening this defensive mechanism through the GNN’s message passing structure.
Enhanced Message Passing. The computed attention weights modulate the message passing in the GNN architecture. For a node $v$, the attention-weighted messages from its neighbors are aggregated as follows:
$\mathbf{h}_{v}^{\ell}\!=\!M(\{\widetilde{A}^{\ell-1}[u,v]\cdot\mathbf{h}_{u}^{\ell-1}\!\mid\!u\!\in\mathcal{N}(v)\}),$ (7)
@@
Unlike NSGT (Qiao & Pang, 2024), which removes edges based on a binary homophily threshold, NSAW introduces a learnable and continuous mechanism to modulate edge importance. This makes NSAW both more flexible and more robust: it preserves useful structure while suppressing anomalous influences during message passing. Instead of pruning edges, NSAW learns to softly attenuate misleading connections, especially those from anomalous nodes, making it more general and data-driven for cross-domain GAD scenarios.
@@ 4.4 Class-aware Regularization
$\mathcal{L}_{s}=\sum_{i\in\mathcal{V}}\left(\frac{\sum_{j\in\mathcal{N}_{i}^{-}}w_{j}\cdot\text{sim}(\mathbf{h}_{i},\mathbf{h}_{j})}{|\mathcal{N}_{i}^{-}|}\right),$ (8)
@@
where $\alpha>1$ is the weighting factor for anomalous nodes. We set $\alpha$ inversely proportional to the class ratio This regularization can only be applied during source training, where we have access to node labels, while target training relies solely on the homophily-based loss due to its unsupervised nature.
@@ 4.5 Test-Time Training
We can now describe our framework for anomaly detection with domain adaptation (GADT3), summarized in Figure 2. Following the test-time training paradigm, GADT3 is based on two tasks, the main (supervised) task and the auxiliary self-supervised learning (SSL) task. In the training phase, GADT3 is trained based on both tasks using a joint loss minimized based on the source dataset. Subsequently, the test-time training phase leverages only the SSL task using the target dataset, which is assumed to be unlabeled.
Training phase. During training, GADT3 applies the source graph $\mathcal{G}_{s}$ to learn an encoder $\theta_{s}$, a decoder $\theta_{m}$, and a predictor $\theta_{\text{pred}}$ by minimizing a loss function $\mathcal{L}_{\text{train}}$ that is a combination of a supervised loss $\mathcal{L}_{\text{sup}}$, for which we apply cross-entropy, and a self-supervised loss $\mathcal{L}_{\text{self}}$, for which we apply the homophily-based affinity scores (see Sec. 4.2):
$\mathcal{L}_{\text{sup}}=\sum_{v\in\mathcal{V}_{s}}y_{v}\log{\hat{y}_{v}}+\lambda_{s}\mathcal{L}_{s},$ (10)
$\mathcal{L}_{\text{self}}=-\sum_{v\in\mathcal{V}}\left(\frac{1}{|N(v)|}\sum_{u\in N(v)}\mathrm{sim}(\mathbf{h}_{v}^{L},\mathbf{h}_{u}^{L})\right)+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (11)
$\mathcal{L}_{\text{train}}=\sum_{v\in\mathcal{V}_{s}}\Bigl(\mathcal{L}_{\text{sup}}(v,y;\theta_{m},\theta_{s},\theta_{\text{pred}})+\lambda\mathcal{L}_{\text{self}}(v;\theta_{m},\theta_{s})\Bigr),$ (12)
@@
We describe the source encoder $\theta_{s}$ (projection) and decoder $\theta_{m}$ (GNN) in Section 4.1. The predictor $\theta_{\text{pred}}$ is a multi-layer perceptron that maps node embeddings to label predictions.
@@
$\mathcal{L}_{\text{TTT}}=\sum_{v\in\mathcal{V}_{t}}\mathcal{L}_{\text{self}}(v;\theta_{t},\theta_{m}),$ (13)
@@
where $\mathcal{L}_{\text{self}}$ is from Eqn. 11 and the decoder $\theta_{m}$ is reused from the training phase to embed target nodes.
Inference: Anomaly detection in the target domain is performed using the adapted encoder $\theta_{t}$ optimized during test-time training (TTT) based solely on the target graph. The decoder $\theta_{m}$, trained on the source domain, is reused without modification. The encoder $\theta_{t}$ maps target node features into a shared embedding space, and $\theta_{m}$ further refines these representations using the target graph. We do not use the prediction head $\theta_{\text{pred}}$ at inference time. Instead, we compute homophily scores in the embedding space generated by $\theta_{m}$, rank the nodes based on these scores, and directly evaluate performance using AUROC and AUPRC.
Although $\theta_{\text{pred}}$ is not applied during inference, we observe the embeddings learned by the adapted target encoder $\theta_{t}$ remain well-aligned with the source decision boundary learned during training. This is shown in Figure 4, where we show that source and target embeddings share similar separation patterns, indicating that the shared representation space preserves discriminative structure even without explicit prediction. While $\theta_{\text{pred}}$ is trained using source labels, it can technically be reused during inference. However, GADT3 relies on homophily-based node ranking. See Appendix A.6 for an alternative approach.